# Curve fitting on a quantum annealer for an advanced navigation method

### v1

@@ IV.1 General formalism
More formally speaking, we consider a deterministic Markov decision process that starts at time $t=0$ and terminates at $t=T$, where the time evolution is discrete in terms of integer time steps: $t=0,\dotsc,T$. The process is characterized by sets $X_{t}\subset\mathbb{R}^{k}$ and $U_{t}\subset\mathbb{R}^{l}$, [^3] In general, the actions depend on the state $x_{t}\in X_{t}$ of the process, i.e., $U_{t}=U_{t}(x_{t})$. However, in terms of notation, we omit this dependency for the the sake of brevity. where $k,l\in\mathbb{N}$, that contain all states the process can be in and all actions that can be performed in order to change the current state, respectively, at time step $t$.
@@
where $D\colon X_{T}\to\mathbb{R}$ encodes terminal costs that occur at the end of the process. The value function is a measure for how cost-effectively the decision process has been if started at time step $t$ with initial state $x_{t}$ and running until the end by virtue of a given sequence of actions $u_{t},\dotsc,u_{T-1}$. The states $x_{t+1},\dotsc,x_{T}$ in Eq. 33 are generated according to the temporal evolution defined by the process’ dynamics 32.
@@
$V_{t}^{\ast}(x_{t})=\min_{\begin{subarray}{c}u_{s}\in U_{s}\\ s=t,\dotsc,T-1\end{subarray}}V_{t}(x_{t},u_{t},\dotsc,u_{T-1})\,.$ (34)
@@ IV.2 Toward an optimized speed profile
  • the trajectory of the vessel is fixed and covers a certain distance of $\ell$ nautical miles;
  • $X_{t}$ contains all feasible positions of the vessel on the trajectory (i.e., the distances traveled since departure at time $t=0$) and environmental aspects like weather, currents, and forecasts at time $t$;
  • $U_{t}$ contains all feasible speeds (through water) of the vessel at time $t$;
  • $C_{t}(x_{t},u_{t})$ are the fuel costs that arise if the vessel moves with speed $u_{t}$ while being at position $x_{t}$ at time $t$;
@@ IV.3 Proof of principle
As a proof of principle, in the following we investigate a simple implementation of what we described above in the previous subsection. We set
@@
Upon inserting these definitions into Eq. 33 and writing the final state as $x_{T}=x_{0}+\sum_{t=0}^{T-1}(u_{t}-w(x_{t}))$, which follows directly from Eq. 39, the value function for the whole process reads
$V_{0}(x_{0},u_{0},\dotsc,u_{T-1})$ $=\frac{1}{v_{\textup{max}}^{2}}\sum_{t=0}^{T-1}(u_{t}-w(x_{t}))^{2}$
$\phantom{=\;}+\frac{\alpha}{\ell}\biggl{[}\ell-x_{0}-\sum_{t=0}^{T-1}(u_{t}-w(x_{t}))\biggr{]}^{2}$
@@
Without loss of generality (and for numerical convenience), we assume that the sets of states and actions $X_{t}$ and $U_{t}$, respectively, are the same for every time step, i.e., independent of $t$. It is then straightforward to find the optimal actions, which are the roots of $\nabla_{u}V_{0}$ (for fixed $x_{0}$). Regarding the gradient of the value function, we find
$\frac{\partial V_{0}}{\partial u_{t}}$ $=\frac{2u_{t}}{v_{\textup{max}}^{2}}-\frac{2\alpha}{\ell^{2}}\biggl{[}\ell-x_{0}-\sum_{s=0}^{T-1}(u_{s}-w(x_{s}))\biggr{]}\,.$ (42)
Now, we shall solve this optimization problem by numerically solving the Bellman equation 35 and subsequently using Eq. 36 to find the optimal actions. For the sake of simplicity, we set $w(x_{t})=0$ for all $x_{t}\in X_{t}$ and all $t$, and consider a voyage with $\ell=100$, $v_{\textup{max}}=50$, and $T=4$. Neglecting phases of (de)acceleration and without currents, the optimal solution is obviously $u_{0}^{\ast}=\dotso=u_{3}^{\ast}=25$, i.e., moving (trivially) with constant velocity. Approaching this problem in terms of the value function 40, we find from $\nabla_{u}V_{0}=0$ (see Eq. 42) the analytical solution
$u_{t}^{\ast}=\frac{\ell}{T+\ell^{2}/(\alpha v_{\textup{max}}^{2})}$ (43)
for all $t=0,\dotsc,3$. As expected, this analytical solution describes a constant time-independent velocity, which consistently coincides with the optimal solution in the limit $\alpha\to\infty$, i.e., for infinite terminal costs. This analytical solution (with finite $\alpha$) shall serve for comparison with results from solving the Bellman equation numerically.
In the course of that, the traditional/naive approach discretizes the state space. Hence, after the value iteration, the resulting optimal value function $V_{t}^{\ast}$ is defined only on that discrete state space grid. Consequently, this poses constraints on the possible discretizations of the action space for the policy iteration 36 because for a given $x_{t}$, $u_{t}$ must be chosen such that $x_{t+1}=\Gamma_{t}(x_{t},u_{t})$ lies on the state space grid, too, or the off-grid result $\Gamma_{t}(x_{t},u_{t})$ has to be “snapped” onto the nearest grid point. Both approaches induce a discretization error that manifests in the final (pseudo-)optimal policy. The in that way obtained optimal policy most likely differs from the true, globally-optimal solution—it is only optimal with respect to the discretization of the state space. In our case, this would result in a pseudo-optimal speed profile different from the analyical solution 43.
@@
$\hat{V}_{t}^{\ast}(x_{t})=\sum_{j=0}^{m-1}c_{j}^{(t)}\phi_{j}(x_{t})$ (44)
with $m\in\mathbb{N}$, and $\phi_{j}\colon X_{t}\to\mathbb{R}$ denote suitable standardized (basis) functions. Here, we used the circumflex notation to indicate that the optimal value function is approximated, and the superscript of the expansion coefficients emphasizes that there is a separate approximation for each time step.
The value iteration takes the same form as before (see Section IV.1) but with the subtle yet important difference, that the optimal value function is approximated at every time step. Starting from the end of the horizon and moving backward in time, the approximation at time step $t$ is then used for $t-1$. We start at the finite horizon with the boundary condition $\hat{V}_{T}^{\ast}(x_{T})=D(x_{T})$ and then proceed backward with
@@
for $t=T-1,\dotsc,0$. Here, it is worth to empasize that at every time step, this value iteration uses only a subset of the state space due to the involved approximation, which allows for an efficient treatment of large state spaces. In the literature, this idea is also known as fitted value iteration [10]. Eventually, the resulting sequence $\hat{V}_{T}^{\ast},\dotsc,\hat{V}_{0}^{\ast}$ of optimal approximated value functions is used in the usual policy iteration 36. Since these optimal value functions are now available for (in principle) arbitrary state space variables thanks to the continuous (yet fitted) nature of Eq. 44, the restrictions for the discretization of the action space are greatly relaxed.
@@
Finally, we would like to present our numercial results. We solve the Bellman equation with its ingredients given in Eqs. 37, 38, 39 and 40. Again, we neglect currents and set $\ell=100$, $v_{\textup{max}}=50$, $T=4$, and $\alpha=1000$. The analytical solution according to Eq. 43 is therefore given by $u_{0}^{\ast}=\dotso=u_{3}^{\ast}=24.975$, which is the optimal solution for our choice of $\alpha$. The Bellman equation is solved grid-based as well as with approximating the optimal value function in the sense of a fitted value iteration. For the latter, the approximation task—which is in this case nothing but curve fitting—is accomplished classically as well as using the QUBO formulation of Section II.2 solved on a quantum annealer. [^5] Again, the Advantage quantum processing unit with 32 runs and 128 reads per run. We employ triangular functions for the expansion 44, $\phi_{j}=\Lambda_{j}$ (see Eq. 4), with $m=5$ and $L=5$, $U=1$ for the QUBO.
@@
Analytical $24.975$ $24.975$ $24.975$ $24.975$ $1.999001$
Grid $24.747$ $24.243$ $25.253$ $25.654$ $1.999453$
Classical $25$ $25$ $25$ $24.90$ $1.999004$
Quantum $25$ $25$ $25$ $24.90$ $1.999004$
Table 3: Optimal policies and their respective total costs for different solution strategies with penalty parameter $\alpha=1000$.
Figure 5: Approximated optimal value functions for each time step with the underlying approximation task of curve fitting for the fitted value iteration accomplished classically (left) and by virtue of quantum annealing.
In Tab. 3, we display our results for the optimal policies obtain by means of the different solution strategies. We find that the grid-based solution is indeed less optimal than the analytical one, which is caused, as explained earlier, by the discretization and forcing off-grid values onto on-grid ones. On the other hand, the solutions using a fitted value iteration are nearly identical with the analytical solution, which is the optimal one for $\alpha=1000$. Moreover, using a QUBO solved on a quantum annealer for the curve fitting in course of the value function approximation yields the same optimal policy than carrying out the fitting classically. Also, on the level of the individual optimal expansion coefficients $c_{j}^{\ast}$ at each nontrivial time step, shown in Tab. 4, we find that quantum annealing delivers, within limits, promising and practically useable results. The difference compared to the classically obtained coefficients is around/below ten percent. For the sake of a visual grasp, in the left (right) diagram of Fig. 5 we display the approximated optimal value function at each time step with the underlying curve fitting done classically (with quantum annealing).
$t=3$ $c_{0}^{\ast}$ $c_{1}^{\ast}$ $c_{2}^{\ast}$ $c_{3}^{\ast}$ $c_{4}^{\ast}$
Classical $3.826$ $4.370$ $2.601$ $1.025$ $1.047$
Quantum $3.936$ $3.969$ $2.781$ $0.969$ $1.125$
$\varepsilon_{\textup{rel}}$ [%] $2.88$ $9.18$ $6.92$ $5.46$ $7.45$
$t=2$ $c_{0}^{\ast}$ $c_{1}^{\ast}$ $c_{2}^{\ast}$ $c_{3}^{\ast}$ $c_{4}^{\ast}$
Classical $3.608$ $2.009$ $1.228$ $1.011$ $1.022$
Quantum $3.875$ $1.969$ $1.156$ $1.031$ $1.031$
$\varepsilon_{\textup{rel}}$ [%] $7.40$ $1.99$ $5.86$ $1.98$ $0.88$
$t=1$ $c_{0}^{\ast}$ $c_{1}^{\ast}$ $c_{2}^{\ast}$ $c_{3}^{\ast}$ $c_{4}^{\ast}$
Classical $2.164$ $1.445$ $1.164$ $1.002$ $1.015$
Quantum $2.000$ $1.344$ $1.156$ $1.031$ $0.938$
$\varepsilon_{\textup{rel}}$ [%] $7.58$ $6.99$ $0.69$ $2.89$ $7.59$
$t=0$ $c_{0}^{\ast}$ $c_{1}^{\ast}$ $c_{2}^{\ast}$ $c_{3}^{\ast}$ $c_{4}^{\ast}$
Classical $1.675$ $1.353$ $1.125$ $0.997$ $1.011$
Quantum $1.563$ $1.344$ $1.156$ $1.063$ $0.969$
$\varepsilon_{\textup{rel}}$ [%] $6.69$ $0.67$ $2.76$ $6.62$ $4.15$
Table 4: Expansion coefficients of the approximated optimal value function obtained classically and via quantum annealing.

### v2

@@ IV.1 General formalism
More formally speaking, we consider a deterministic Markov decision process that starts at time $t=0$ and terminates at $t=T$, where the time evolution is discrete in terms of integer time steps: $t=0,\dotsc,T$. The process is characterized by sets $X_{t}\subset\mathbb{R}^{k}$ and $U_{t}\subset\mathbb{R}^{l}$, where $k,l\in\mathbb{N}$, that contain all states the process can be in and all actions that can be performed in order to change the current state, respectively, at time step $t$ [^4] In general, the actions depend on the state $x_{t}\in X_{t}$ of the process, i.e., $U_{t}=U_{t}(x_{t})$. However, in terms of notation, we omit this dependency for the the sake of brevity.
@@
where $D\colon X_{T}\to\mathbb{R}$ encodes terminal costs that occur at the end of the process. The value function is a measure for how cost-effectively the decision process has been if started at time step $t$ with initial state $x_{t}$ and running until the end by virtue of a given sequence of actions $u_{t},\dotsc,u_{T-1}$. The states $x_{t+1},\dotsc,x_{T}$ in Eq. (34) are generated according to the temporal evolution defined by the process’ dynamics 33.
@@
$V_{t}^{\ast}(x_{t})=\min_{\begin{subarray}{c}u_{s}\in U_{s}\\ s=t,\dotsc,T-1\end{subarray}}V_{t}(x_{t},u_{s},\dotsc,u_{T-1})\,.$ (35)
@@ IV.2 Toward an optimized speed profile
  • The trajectory of the vessel is fixed and covers a certain distance of $\ell$ nautical miles.
  • $X_{t}$ contains all feasible positions of the vessel on the trajectory (i.e., the distances traveled since departure at time $t=0$) and environmental aspects like weather, currents, and forecasts at time $t$.
  • $U_{t}$ contains all feasible speeds (through water) of the vessel at time $t$.
  • $C_{t}(x_{t},u_{t})$ are the fuel costs that arise if the vessel moves with speed $u_{t}$ while being at position $x_{t}$ at time $t$.
@@ IV.3 Proof of principle
For the purpose of a proof of principle, we choose a linear-quadratic optimal control problem. The system’s dynamics are linear in both the state and action variables, while the cost functional comprises both running and terminal costs, each formulated as quadratic forms. This specific structure offers several analytical and computational advantages. First, the convexity of the cost functional and the linearity of the system’s dynamics ensure that the resulting optimization problem is convex. This implies the existence of an unique global optimum. Second, the problem admits an analytical solution via the Riccati differential equation, enabling the derivation of the optimal control law in the form of a linear state feedback.
In the following, we investigate a simple implementation of what we described above in the previous subsection, namely the problem of just-in-time arrival. We set
@@
Upon inserting these definitions into Eq. 34 and writing the final state as $x_{T}=x_{s}+\sum_{t=s}^{T-1}(u_{t}-w(x_{t}))$ as a sum of actions starting at time $s$ in state $x_{s}$, which follows directly from Eq. 40, the value function for the whole process reads
$V_{s}(x_{s},u_{s},\dotsc,$ $\,u_{T-1})=\frac{1}{v_{\textup{max}}^{2}}\sum_{t=s}^{T-1}(u_{t}-w(x_{t}))^{2}$
$\phantom{=\;}+\alpha\biggl{[}\ell-\frac{1}{\ell}\biggl{(}x_{s}-\sum_{t=s}^{T-1}(u_{t}-w(x_{t}))\biggr{)}\biggr{]}^{2}$
@@
Without loss of generality (and for numerical convenience), we assume that the sets of states and actions $X_{t}$ and $U_{t}$, respectively, are the same for every time step, i.e., independent of $t$. It is then straightforward to find the optimal actions, which are the roots of $\nabla_{u}V_{s}$ (for fixed $x_{s}$). Regarding the gradient of the value function, we find
$\frac{\partial V_{0}}{\partial u_{t}}=\frac{2u_{t}}{v_{\textup{max}}^{2}}-\frac{2\alpha}{\ell^{2}}\biggl{[}\ell-x_{s}-\sum_{s^{\prime}=s}^{T-1}(u_{s^{\prime}}-w(x_{s^{\prime}}))\biggr{]}\,.$ (43)
Now, we shall solve this optimization problem by numerically solving the Bellman equation (36) and subsequently using Eq. 37 to find the optimal actions. For the sake of simplicity, we set $w(x_{t})=0$ for all $x_{t}\in X_{t}$ and all $t$, and consider a voyage with $\ell=100$, $v_{\textup{max}}=50$, and $T=4$. Neglecting phases of (de)acceleration and without currents, the optimal solution is obviously $u_{0}^{\ast}=\dotso=u_{3}^{\ast}=25$, i.e., moving (trivially) with constant velocity. Approaching this problem in terms of the value function 41, we find from Eq. (43) the analytical solution
$u_{t}^{\ast}=\frac{\alpha v_{\textup{max}}^{2}(\ell-x_{s})}{T-s+\ell^{2}}$ (44)
for all $t=s,\dotsc,T$. As expected, this analytical solution describes a constant time-independent velocity, which consistently coincides with the optimal solution in the limit $\alpha\to\infty$, i.e., for infinite terminal costs. This analytical solution (with finite $\alpha$) shall serve for comparison with results from solving the Bellman equation numerically.
In the course of that, the traditional/naive approach discretizes the state space. Hence, after the value iteration, the resulting optimal value function $V_{t}^{\ast}$ is defined only on that discrete state space grid. Consequently, this poses constraints on the possible discretizations of the action space for the policy iteration 37 because for a given $x_{t}$, $u_{t}$ must be chosen such that $x_{t+1}=\Gamma_{t}(x_{t},u_{t})$ lies on the state space grid, too, or the off-grid result $\Gamma_{t}(x_{t},u_{t})$ has to be “snapped” onto the nearest grid point. Both approaches induce a discretization error that manifests in the final (pseudo-)optimal policy. The in that way obtained optimal policy most likely differs from the true, globally-optimal solution—it is only optimal with respect to the discretization of the state space. In our case, this would result in a pseudo-optimal speed profile different from the analytical solution 44.
@@
$\hat{V}_{t}^{\ast}(x_{t};c)=\sum_{j=0}^{m-1}c_{j}^{(t)}\phi_{j}(x_{t})$ (45)
with $m\in\mathbb{N}$, and $\phi_{j}\colon X_{t}\to\mathbb{R}$ denote suitable standardized (basis) functions. Here, we used the circumflex notation to indicate that the optimal value function is approximated, and the superscript of the expansion coefficients emphasizes that there is a separate approximation for each time step. For the sake of brevity, we omit the dependence of $\hat{V}_{t}^{\ast}$ on the coefficient vector $c$ for the remainder of this section. The value iteration takes the same form as before (see Section IV.1) but with the subtle yet important difference that the optimal value function is approximated at every time step. Starting from the end of the horizon and moving backward in time, the approximation at time step $t$ is then used for $t-1$. We start at the finite horizon $t=T$ with the boundary condition $\hat{V}_{T}^{\ast}(x_{T})=D(x_{T})$ and then proceed backward with
@@
for $t=T-1,\dotsc,0$. Here, it is worth to emphasize that at every time step, this value iteration uses only a subset of the state space due to the involved approximation, which allows for an efficient treatment of large state spaces. In the literature, this idea is also known as fitted value iteration [10]. Eventually, the resulting sequence $\hat{V}_{T}^{\ast},\dotsc,\hat{V}_{0}^{\ast}$ of optimal approximated value functions is used in the usual policy iteration 37. Since these optimal value functions are now available for (in principle) arbitrary state space variables thanks to the continuous (yet fitted) nature of Eq. 45, the restrictions for the discretization of the action space are greatly relaxed.
@@
Now, we would like to present our numerical results. We solve the Bellman equation with its ingredients given in Eqs. (38)–(41). Again, we neglect currents and set $\ell=100$, $v_{\textup{max}}=50$, and $T=4$. The analytical solution according to Eq. 44 is therefore given by
$u_{0}^{\ast}=\dotso=u_{3}^{\ast}=\frac{25}{1+1/\alpha}\,,$ (47)
which is the optimal solution for an arbitrary $\alpha<\infty$. The Bellman equation is solved grid-based as well as with approximating the optimal value function in the sense of a fitted value iteration. For the latter, the approximation task—which is nothing but curve fitting—is accomplished classically as well as using the QUBO formulation of Section II.2 solved on a quantum annealer. [^6] Again, the Advantage2 system but with 50 runs, 256 reads per run, and an annealing time of 1000 microseconds. We employ triangular functions for the expansion 45, $\phi_{j}=\Lambda_{j}$ (see Eq. 4), with $m=9$, $d=9$, and $p=8$ for the QUBO.
Figure 5: Visualization of the value function e approximation using different solution strategies (upper row) and their corresponding mean squared errors (MSE) (lower row). The left column shows the results for time slice $t=3$, while the right column presents the same analysis for the first time slice $t=0$. The upper panels depict the approximated value functions, whereas the lower panels show the associated approximation errors. Here and in the following figures, we use the MSE in contrast to the RMSE as defined in Eq. 26.
In the following bullet points, we summarize our solution strategies for the Bellman equation. The last three methods use a fitted value iteration as described above, where the approximation task is solved classically by matrix inversion or via a QUBO using a tabu search as well as quantum annealing. In particular, we have:
  • Analytic: Approaching the problem in terms of the value function 41, we find from $\nabla_{u}V_{s}=0$ (see Eq. 43) the analytical solution.
  • Grid-based: This baseline method avoids functional approximation entirely. Instead, the state space is discretized on a equidistant grid and estimation of the remaining costs-to-go is performed directly on grid points. It represents a brute-force approach, useful for benchmarking, but it is computationally expensive in high-dimensional state spaces.
  • Classically (matrix inversion): This method relies on classical linear algebra techniques. The system of equations resulting from the linear basis function ansatz is solved using Gaussian elimination. It serves as a deterministic baseline and provides exact coefficients.
  • Tabu search: The QUBO formulation of the approximation problem is solved by a tabu search, a metaheuristic method well-suited for combinatorial landscapes. It offers a flexible alternative when exact methods are computationally unfeasible.
  • Quantum annealing: The QUBO formulation of the approximation problem is submitted to a D-Wave quantum annealer. This explores the potential of currently available quantum annealing devices.
In Tab. 5, we display our results for the optimal policies obtained by means of the different solution strategies. We have an example where the grid-based solution is indeed less optimal than the results that employ the fitted value iteration with the approximation solved classically by matrix inversion and QUBO-based via a tabu search. This is caused, as explained earlier, by the discretization and forcing off-grid values onto on-grid ones. Moreover, using a QUBO solved on a quantum annealer for the curve fitting in course of the fitted value iteration yields a worse optimal policy.
Table 5: Optimal policies and their respective total costs for different solution strategies with penalty parameter $\alpha=100$, $m=9$, and $d=9$.
@@
analytic $24.752$ $2.752$ $24.752$ $24.752$ $1.990099$
grid $23.470$ $24.490$ $24.490$ $26.492$ $1.992070$
inverse $25$ $25$ $25$ $24.039$ $1.990384$
tabu $25$ $25$ $25$ $24.039$ $1.990384$
qa $12.5$ $12.5$ $12.5$ $50$ $2.597069$
Figure 6: MSE of the reconstructed solution $\hat{V}^{\ast}$ compared to the reference solution $V^{\ast}$. They carry no time index because we averaged over all time slices. Left: Convergence behavior of the MSE with respect to the number of discretization points $n$ of the state space for different regularization parameters $\alpha\in\{10,100,1000\}$. Right: Error decay with respect to the number of basis functions $m$ for fixed discretization $n=50$ and varying $\alpha$. The horizontal dashed-dotted line indicates the target error threshold, where the approximation solution with respect to the number of basis functions $m$ starts to be more accurate than the grid solution.
Figure 7: Error comparison of the two solution approaches, quantum annealing and tabu search, at different precision levels ($d=7$ and $d=9$), as a function of the number of basis functions $m$. The left panel shows the results for the first time slice in the fitted value iteration process, while the right panel displays the average performance over all time slices.
It should be noted that the choice of $m$ and $d$ results in an $81$-dimensional QUBO problem. The annealer does not solve this problem size accurately enough, which is consistent with the findings in Section III. Fig. 5 shows that the solution of the quantum annealing already exhibits a significant error at the very first time slice, preventing the method from reaching the optimal solution. Due to error propagation, this issue is further exacerbated in the subsequent time slices. The reasons lie in the size of the search space and the specific energy landscape formed by the binary representation of the coefficients.
The example calculated above does not provide a clear picture of the behavior if one varies the number of triangular functions $m$. Therefore, we varied the latter and analyzed the resulting errors. As shown in Fig. 6, increasing the number of basis functions leads to improved approximation results. This holds also for other values of $\alpha$. At this point, all calculations shown in Fig. 6 are classical solutions of the QUBO formulation of the approximation task for $\hat{V}_{t}^{\ast}$ using Gaussian elimination. Basically, this procedure corresponds to computing a matrix inverse. The resulting solution can be interpreted as the best possible outcome within the span of the chosen basis. In a following figure, we further investigate how different solvers approximate this best solution. In summary, increasing the number of basis functions effectively reduces the approximation error. The remaining challenge is to ensure that the solvers are capable of achieving this improved level of solution quality.
Our analysis shows in Fig. 7 that as the number of basis functions—and thus the problem size—increases, the solutions generally deviate further from the best possible solution, referred to as the inverse solution (indicated in blue). This trend holds for the tabu search but is especially pronounced in the case of quantum annealing. Starting from six basis functions, the problem size becomes large enough for the approximation error to increase significantly. For smaller problem sizes, the quantum annealing yields approximation results that are comparable to those obtained with the classical solution or the tabu search. However, beyond a certain problem size, this similarity breaks down as clearly illustrated in Fig. 7.
IV.4 Further discussion
The proposed QUBO-based approach relates to a broader class of modern methods developed to approximate solutions of high-dimensional HJB equations, which typically arise in optimal control and stochastic decision problems. More specifically, we investigate a problem in the context of approximate dynamic programming with a bounded horizon. We do not consider broader reinforcement learning techniques that incorporate policy optimization or explicit exploration-exploitation mechanisms. Furthermore, no explicit error-control strategies are employed, as our focus lies solely on the approximation quality of the value function using the QUBO approach. A central challenge in solving such equations numerically is the curse of dimensionality, i.e., the exponential growth of computational complexity with respect to the state and action space dimensions. The QUBO-based approach can be extended to multidimensional settings by employing, for example, a tensor product of basis functions.
Recent literature has addressed this challenge using learning-based strategies, most notably through the approximation of the value function or the HJB operator via deep neural networks. Notable examples include the Deep Galerkin Method [68] and physics-informed neural networks [69]. Both utilize neural networks as function approximators trained on residuals of partial differential equations. These methods are well-suited for high-dimensional spaces by leveraging the expressive capacity and scalability of deep architectures.
In contrast to these continuous, gradient-based approaches, our method introduces a discrete, combinatorial framework for function approximation based on a QUBO. This discrete optimization perspective can be interpreted as a learning approach that bypasses gradient descent and instead exploits a new perspective on value function approximation, particularly in problem settings that can be nonconvex or nondifferentiable. In such cases, the resulting value function may exhibit sharp transitions or multiple local extrema. When approximating such functions using triangular functions on a discretized grid, the associated QUBO landscape can become irregular, potentially containing many local minima. This may lead to instabilities in the QUBO optimization and reduced approximation accuracy. However, this is not an intrinsic limitation of the QUBO approach itself, but rather a structural artifact of the discretization and basis choice. In fact, quantum annealing metaheuristics, such as those employed by D-Wave, offer the potential to escape local minima via quantum tunneling, which may be advantageous in overcoming the challenges of nonconvexity.
Therefore, while nonconvexity and nondifferentiability may degrade the approximation landscape, the global nature of the QUBO formulation together with carefully chosen basis functions preserves the possibility of capturing the correct global structure of the value function.